OpenStudy (anonymous):

evaluate 7^-1 - 7^0 (show working)?

OpenStudy (anonymous):

7^(-1) is the same as 1/7 anything raised to the power of zero is 1 1/7-1 1/7-7/7 -6/7

OpenStudy (anonymous):

7^-1 = 1/7 (This is by definition, put number with negative exponent (-1 in this case) below. Work: 7^-1 = 1/(7^1) , any number to the 1 power is itself) 7 ^ 0 = 1 (This is also by definition, any number to the power of 0 is equal to 1) Therefore, 7^-1 - 7^0 = 1/7 - 1 = 1/7 - 7/7 = -6/7